# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v2

@@ 5.1 Analysis
In this section we compare the robustness of REVS with FT-L, the second-best performing method, under different hyperparameter settings. Figure 4 shows that REVS Pareto-dominates FT-L in all but one experiment.
@@
REVS consistently outperforms FT-L as we increase the size of the candidate token set $C$ (Figures 4 and 4). While for the Email dataset both methods exhibit similar trends, REVS maintains superior performance for each size of $C$. For the SSN dataset, REVS demonstrates near-perfect efficacy across all candidate token set sizes. On the Email dataset both methods exhibit similar trends, but REVS maintains superior performance for each $C$. On the SSN dataset, REVS reports near-perfect efficacy across all candidate token set sizes.
Efficacy/specificity trade–off.
While both REVS and FT-L are affected by this trade-off (Figures 4 and 4), REVS fares better, robustly maintaining high specificity when increasing edit strength. Notably, on SSN, there is almost no trade-off, with REVS achieving near-perfect specificity and efficacy.
@@
Figure 4 shows that REVS maintains a stable Unlearning Score across different target edit numbers while FT-L performs well only when the number of edits matches its tuned hyperparameters, leading to poor performance for other target quantities. Figure 4 shows a similar overall downward trend in Unlearning Score as the number of edits increases for both REVS and FT-L. However, REVS consistently maintains superior performance across all edit quantities.
Robustness.
A key strength of REVS is its robustness across different datasets. Using the same hyperparameter configuration for both SSN and Emails, REVS outperforms other baseline methods that were individually tuned for each specific dataset (Section B.2).
@@ 6 Related Work
Privacy concerns have given way to development of methods preventing regurgitation of PII. Differential privacy (DP; Dwork et al., 2014; Ramaswamy et al., 2020; Hoory et al., 2021; Li et al., 2022) and dataset scrubbing through deduplication (Lee et al., 2022; Kandpal et al., 2022; Carlini et al., 2023a) prevent sensitive information from ever being stored in the model. However, using DP-SGD frequently results in worse generative models (Anil et al., 2022), and LMs are able to memorize even singleton training instances (Debenedetti et al., 2024). Scrubbing approaches are also expensive, requiring retraining for each new identified PII leak. To alleviate high cost incurred by retraining models from scratch, post-hoc model unlearning techniques (Liu et al., 2024) such as gradient ascent (Bourtoule et al., 2021; Neel et al., 2021; Eldan and Russinovich, 2023), reinforcement learning through human feedback (RLHF; Ouyang et al., 2022), and model editing (Cao et al., 2021; Meng et al., 2022, 2023; Wu et al., 2023) were devised. Most similar to ours is the work of Wu et al. (2023), who identify and zero out neurons responsible for memorized sensitive information. REVS diverges in key aspects, significantly outperforming naive zeroing by modifying selected neurons to demote relevant tokens in the vocabulary space, thus removing the tendency to generate sensitive information while preserving model integrity (see Section E.2).
@@
Membership inference attacks (Dwork et al., 2006; Shokri et al., 2017) try to verify whether a specific datum was in a model’s training set, while attribute inference attacks (Fredrikson et al., 2014) retrieve information on some attributes of the training instances. Data extraction attacks aim to retrieve full training instances. Generative LMs are especially susceptible to such attacks given their propensity to memorize training data (Carlini et al., 2019, 2023b). Patil et al. (2024) show that despite efforts to scrub sensitive information from models, different attacks can still locate such information in LMs. Black-box attacks retrieve information without access to model internals (Henderson et al., 2018; Lukas et al., 2023; Krishna et al., 2024), while white-box attacks use model parameters and probing (Nostalgebraist, 2020; Geva et al., 2022) to even detect information that has been erased by model editing techniques (Patil et al., 2024). REVS is specifically designed with such attacks in mind and removes sensitive information in a way robust with respect to possible attacks as well as effective in terms of erasing information.
@@ 7 Conclusion
We introduce REVS, a novel unlearning technique which demonstrates superior performance to state-of-the-art baselines in effectively unlearning sensitive information from LMs while preserving model integrity, general capabilities and remaining robust to extraction attacks. We curate two datasets containing sensitive information, one organically memorized by the models during training. Through detailed analyses we show REVS’ robustness to varying candidate token set sizes, number of unlearned targets and the trade-off between efficacy and specificity.
@@
While REVS shows promising results, several areas need further work to resolve following limitations: (1) generalization to unlearning other types of organically memorized sensitive information beyond emails, (2) scalable unlearning of larger amounts of target information, (3) evaluation against more fine-grained extraction attacks, including on specific sub-modules like MLP or Attention outputs, (4) evaluation on datasets in languages other than English.
@@ Appendix A Results
A.1 Full Results of Extraction Resistance on Llama-3-8B
@@
FT-L $64.35{\pm 1.62}$ $55.07{\pm 2.40}$ $92.43{\pm 0.80}$ $56.90{\pm 3.34}$
MEMIT $66.93{\pm 3.66}$ $51.12{\pm 4.85}$ $92.83{\pm 0.23}$ $\underline{69.45{\pm 3.26}}$
Max-Entropy $\underline{73.18{\pm 3.67}}$ $\bm{75.15{\pm 2.44}}$ $\bm{93.20{\pm 0.40}}$ $59.45{\pm 7.11}$
Head-Projection $69.37{\pm 4.08}$ $67.05{\pm 2.85}$ $92.32{\pm 0.23}$ $57.57{\pm 7.15}$
RMU $26.10{\pm 15.04}$ $13.80{\pm 10.05}$ $91.80{\pm 0.96}$ $51.53{\pm 3.73}$
NPO-KL $39.80{\pm 6.95}$ $22.90{\pm 5.47}$ $91.75{\pm 0.67}$ $49.82{\pm 5.87}$
REVS (ours) $\bm{79.78{\pm 1.74}}$ $\underline{74.20{\pm 2.33}}$ $\underline{93.13{\pm 0.68}}$ $\bm{74.83{\pm 2.65}}$
@@ A.2 Results on GPT-J-6B
Even after few-shot learning, the baseline unedited model performed near random chance on MMLU, achieving scores of $27$ for Emails and $26$ for SSN, where $25$ represents random performance. Consequently, we did not evaluate MMLU scores for GPT-J-6B.
@@ A.3 Analysis
Efficacy-specificity trade-off
@@ B.1 Implementation and Experimental Details
B.2 Hyperparameter Robustness Evaluation
We analyze the performance of REVS for Llam-3-8B using a fixed hyperparameter configuration for both datasets in Table 5. The results demonstrate REVS’s robust performance on both the SSN and Emails datasets when using the same set of hyperparameters. For the SSN dataset, REVS significantly outperforms FT-L across most metrics, showcasing its effectiveness in handling sensitive information. In the Emails dataset, our method achieves notably high performance, particularly in efficacy, while maintaining competitive scores across other evaluation metrics. These findings underscore the versatility of REVS and its potential for generalized application across different domains and data types.
Dataset Method Unlearning Score $\uparrow$ Efficacy@100 $\uparrow$ Specificity $\uparrow$ Generality@100 $\uparrow$ Resistance Score $\uparrow$
SSN FT-L $36.98{\pm 11.97}$ $63.88{\pm 9.88}$ $24.33{\pm 9.78}$ $50.35{\pm 10.76}$ $82.77{\pm 5.73}$
REVS (same HP) $\bm{70.73{\pm 4.68}}$ $\bm{75.85{\pm 4.97}}$ $\bm{92.67{\pm 1.60}}$ $\bm{54.70{\pm 6.49}}$ $\bm{90.02{\pm 2.23}}$
Emails FT-L $58.20{\pm 2.42}$ $55.07{\pm 2.40}$ $\bm{61.91{\pm 4.19}}$- $64.35{\pm 1.62}$
REVS (same HP) $\bm{62.62{\pm 6.17}}$ $\bm{93.15{\pm 1.91}}$ $47.62{\pm 7.04}$- $\bm{87.07{\pm 1.63}}$
Table 5: Comparison of REVS using the same hyperparameter (HP) configuration across both datasets to the second-best method (FT-L), which employs custom HP configurations for each dataset. Results are shown for Llama-3-8B.
@@ B.3 Editing Neuron
B.4 Qualitative Unlearning Examples
We demonstrate the Unlearning Effectiveness and Model Integrity of REVS through qualitative examples in Tables 6 and 7. Table 6 presents randomly sampled instances from our dataset, showcasing how the unlearning process handles private email addresses. After applying our method, the model generates plausible alternatives to the original private email addresses. Specifically, we observe two primary transformation cases: (1) replacing the original email address with a similar but different email address that maintains the contextual integrity of the original text; (2) replacing the specific email with a generic [email protected] placeholder while preserving the overall coherence and meaning of the surrounding text.
@@ C.1 MEMIT Baseline
For the Llama-3-8B model, we applied the method with specific hyperparameters: on the SSN dataset, we focused on layer $3$ with a loss break of $1\times 10^{-5}$, loss prediction probability coefficient of $100$, learning rate of $0.01$, and $5$ gradient steps with a maximum prompt length of $n_{max\_len}=20$. For the EMAIL dataset, we maintained similar settings, adjusting only the maximum prompt length to $n_{max\_len}=100$. For the GPT-J-6B model, the following hyperparameters were used: on the SSN dataset, we targeted layers $3$-$8$ with a loss break of $0.1$, loss prediction probability coefficient of $20$, learning rate of $0.05$, $25$ gradient steps, and a maximum prompt length of $n_{max\_len}=200$. The EMAIL dataset configuration was similar, with a reduced maximum prompt length of $n_{max\_len}=50$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: velocity learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.5 Head-Projection Baseline
We followed the original implementation with a modified objective to prevent the target from appearing in the top $k=100$ logit distribution elements. For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration used a loss break of $1.9\times 10^{-3}$, loss prediction probability coefficient of $3.08\times 10^{-2}$, learning rate of $1.03\times 10^{-2}$, and $10$ gradient steps. The EMAIL used a loss break of $7.6\times 10^{-4}$, loss prediction probability coefficient of $2.58\times 10^{-1}$, learning rate of $9.19\times 10^{-3}$, and $25$ gradient steps. For the GPT-J-6B model we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset we used a loss break of $6.596\times 10^{-2}$, loss prediction probability coefficient of $3.271$, learning rate of $2.045\times 10^{-1}$, and $25$ gradient steps. The EMAIL dataset configuration used similar layer targeting but modified the hyperparameters, with a loss break of $6\times 10^{-4}$, loss prediction probability coefficient of $6.733\times 10^{-1}$, learning rate of $1.868\times 10^{-1}$, and $5$ gradient steps. In our hyperparameter sweep, we considered the following ranges: velocity learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.6 Max-Entropy Baseline
For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration employed a loss break of $1.99\times 10^{-2}$, loss prediction probability coefficient of $20.5$, learning rate of $7.6\times 10^{-3}$, and $25$ gradient steps. The EMAIL dataset adjusted these parameters, using a loss break of $1.75\times 10^{-3}$, loss prediction probability coefficient of $1.35\times 10^{-2}$, learning rate of $9.2\times 10^{-3}$, and $5$ gradient steps. For the GPT-J-6B model, we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset, we used a loss break of $7.0559\times 10^{-2}$, loss prediction probability coefficient of $25.9$, learning rate of $1.937\times 10^{-1}$, and $5$ gradient steps. The EMAIL dataset used a loss break of $3.634\times 10^{-2}$, loss prediction probability coefficient of $1.907\times 10^{-2}$, learning rate of $6.65\times 10^{-2}$, and $10$ gradient steps. In our hyperparameter sweep, we considered the following ranges: velocity learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ E.1 Details and Examples
In this work, all methods unlearn the same target tokens from the sequence of sensitive information. Our approach strategically focuses on identifying and targeting the rarest $T\subseteq S$ tokens from each sequence $S$, to minimize unnecessary model perturbation while effectively isolating sensitive information. We approximate the rareness of a token by its ID as assigned by the tokenizer, where a higher ID value indicates a rarer token.
[…]
This strategy allows us to remove the most unique, information-dense tokens that are most likely to contain sensitive identifying information, while minimizing unnecessary model perturbation.

### v3

@@ 5.1 Analysis
In this section we compare the robustness of REVS with FT-L, the second-best performing method, under different hyperparameters. Figure 4 shows that REVS Pareto-dominates FT-L in all but one experiment.
@@
REVS consistently outperforms FT-L as we increase the size of the candidate token set $C$ (Figures 4, 4 and 4). While for the Emails and URLs datasets both methods exhibit similar trends, REVS maintains superior performance for each size of $C$. For the SSN dataset, REVS demonstrates near-perfect efficacy across all candidate token set sizes.
Efficacy/Specificity trade-off.
While both REVS and FT-L are affected by this trade-off (Figures 4, 4 and 4), REVS fares better, robustly maintaining high specificity when increasing edit strength. Notably, on SSN, there is almost no trade-off, with REVS achieving near-perfect specificity and efficacy.
@@
On SSN (Figure 4) REVS maintains a stable Unlearning Score across edit counts, while FT-L performs well only at its tuned hyperparameters, degrading elsewhere. Organic data (Figures 4 and 4) show a downward trend in Unlearning Score for both methods as edits increase; however, REVS consistently outperforms FT-L.
@@ 6 Related Work
Privacy concerns have given way to development of methods preventing regurgitation of PII. Differential privacy (DP; Dwork et al., 2014; Ramaswamy et al., 2020; Hoory et al., 2021; Li et al., 2022) and dataset scrubbing through deduplication (Lee et al., 2022; Kandpal et al., 2022; Carlini et al., 2023a) prevent sensitive information from ever being stored in the model. However, using DP-SGD frequently results in worse generative models (Anil et al., 2022), and LMs can memorize even singleton training instances (Debenedetti et al., 2024). Scrubbing approaches are also expensive, requiring retraining for each new identified PII. To alleviate high cost incurred by retraining models, post-hoc model unlearning techniques (Liu et al., 2024) such as gradient ascent (Bourtoule et al., 2021; Neel et al., 2021; Eldan and Russinovich, 2023), reinforcement learning through human feedback (RLHF; Ouyang et al., 2022), and model editing (Cao et al., 2021; Meng et al., 2022, 2023; Wu et al., 2023) were devised. Most similar to ours is the work of Wu et al. (2023), who identify and zero out neurons responsible for memorized sensitive information. REVS diverges in key aspects, significantly outperforming naive zeroing by modifying selected neurons to demote relevant tokens in the vocabulary space, thus removing the tendency to generate sensitive information while preserving model integrity (see Section E.2).
@@
Membership inference attacks (Dwork et al., 2006; Shokri et al., 2017) try to verify whether a specific datum was in a model’s training set, while attribute inference attacks (Fredrikson et al., 2014) retrieve information on attributes of training instances. Data extraction attacks aim to retrieve full training instances. Generative LMs are susceptible to such attacks given their propensity to memorize training data (Carlini et al., 2019, 2023b). Patil et al. (2024) show that despite efforts to scrub sensitive information from models, different attacks can still locate such information in LMs. Black-box attacks retrieve information without access to model internals (Henderson et al., 2018; Lukas et al., 2023; Krishna et al., 2024), while white-box attacks use model parameters and probing (Nostalgebraist, 2020; Geva et al., 2022) to even detect information erased by model editing techniques (Patil et al., 2024). REVS is specifically designed with such attacks in mind, removing sensitive information effectively while maintaining robustness against potential attacks.
@@ 7 Conclusion
We introduce REVS, a novel unlearning method which demonstrates superior performance to state-of-the-art baselines in effectively unlearning sensitive information from LMs while preserving general capabilities and remaining robust to extraction attacks. We curate three datasets containing sensitive information, two naturally memorized by the models. Through detailed analyses we show REVS’ robustness to varying candidate token set sizes, number of unlearned targets and the trade-off between efficacy and specificity.
@@
While REVS shows promising results, several areas require further work: (1) scalability of unlearning to larger amounts of target information, (2) generalization to unlearning other types of PII, particularly numeric, while preserving model integrity, (3) evaluation against more fine-grained extraction attacks, including targeted analysis of MLP and Attention modules, as well as probe-based assessments of residual information, (4) evaluation on datasets in languages other than English.
@@ Appendix A Results
A.1 Perturbation Attack Analysis.
As shown in Tables 2 and 4, the Perturbation Attack (PA) score is relatively low before unlearning, suggesting that the target tokens are already highly ranked in the model’s output given the perturbed prompt. After unlearning, the PA score increases significantly. Notably, for REVS, the PA score doubles post-unlearning, indicating its effectiveness against this attack. Interestingly, on Llama-3-8B for the URL dataset, some methods (e.g., RMU and NPO) exhibit a lower PA score than the unedited model, suggesting that internal memorization persists and becomes more pronounced when the original prompt is used for unlearning with these methods.
A.2 Results on Llama-3-8B
Full Results of Extraction Resistance.
@@ A.2 Results on Llama-3-8B
FT-L $64.77{\pm 3.93}$ $52.98{\pm 4.23}$ $\underline{92.27{\pm 0.58}}$ $60.65{\pm 6.20}$
MEMIT $70.55{\pm 4.00}$ $63.63{\pm 3.50}$ $92.12{\pm 1.10}$ $\underline{62.95{\pm 6.20}}$
Max-Entropy $\bm{72.77{\pm 3.90}}$ $\bm{69.75{\pm 6.30}}$ $\bm{92.47{\pm 1.00}}$ $62.47{\pm 3.60}$
Head-Projection $70.28{\pm 3.70}$ $\underline{64.33{\pm 4.90}}$ $92.10{\pm 0.60}$ $61.57{\pm 4.10}$
RMU $29.82{\pm 8.40}$ $15.08{\pm 5.90}$ $91.05{\pm 0.90}$ $48.23{\pm 4.00}$
NPO-KL $42.05{\pm 3.60}$ $24.27{\pm 3.00}$ $91.17{\pm 0.40}$ $52.90{\pm 4.10}$
REVS (ours) $\underline{71.95{\pm 2.25}}$ $59.65{\pm 3.97}$ $91.85{\pm 0.62}$ $\bm{71.47{\pm 1.65}}$
URLs Unedited $0.00{\pm 0.00}$ $0.00{\pm 0.00}$ $91.00{\pm 0.00}$ $36.00{\pm 0.00}$
FT-L $\underline{56.70{\pm 10.16}}$ $\underline{58.80{\pm 7.37}}$ $\bm{95.58{\pm 0.44}}$ $\underline{41.18{\pm 13.90}}$
MEMIT $45.30{\pm 10.30}$ $34.37{\pm 10.80}$ $92.90{\pm 0.90}$ $39.35{\pm 9.60}$
Max-Entropy $44.93{\pm 7.90}$ $32.88{\pm 7.90}$ $\underline{94.92{\pm 0.50}}$ $39.50{\pm 6.80}$
Head-Projection $37.83{\pm 9.50}$ $26.32{\pm 8.40}$ $93.63{\pm 1.80}$ $33.63{\pm 7.50}$
RMU $16.88{\pm 15.10}$ $12.22{\pm 12.20}$ $92.48{\pm 1.30}$ $13.88{\pm 13.10}$
NPO-KL $16.77{\pm 6.80}$ $10.97{\pm 4.40}$ $91.08{\pm 0.40}$ $13.50{\pm 6.30}$
REVS (ours) $\bm{82.80{\pm 3.94}}$ $\bm{78.22{\pm 6.04}}$ $91.97{\pm 0.15}$ $\bm{79.75{\pm 4.96}}$
@@ A.3 Results on GPT-J-6B
Even after few-shot learning, the baseline unedited model performed near random chance on MMLU, achieving scores of $27$ for Emails and $26$ for SSN, where $25$ represents random performance. Consequently, we did not evaluate MMLU and GSM8K scores for GPT-J-6B.
@@ A.4 Analysis
Efficacy/Specificity trade-off
@@ B.2 Editing Neuron
B.3 Qualitative Unlearning Examples
We demonstrate the Unlearning Effectiveness and Model Integrity of REVS through qualitative examples in Tables 5 and 6. Table 5 presents randomly sampled instances from our dataset, showcasing how the unlearning process handles private email addresses. After applying our method, the model generates plausible alternatives to the original private email addresses. Specifically, we observe two primary transformation cases: (1) replacing the original email address with a similar but different email address that maintains the contextual integrity of the original text; (2) replacing the specific email with a generic [email protected] placeholder while preserving the overall coherence and meaning of the surrounding text.
@@ C.1 MEMIT Baseline
For the Llama-3-8B model, we applied the method with specific hyperparameters: on the SSN dataset, we focused on layer $3$ with a loss break of $1\times 10^{-5}$, loss prediction probability coefficient of $100$, learning rate of $0.01$, and $5$ gradient steps with a maximum prompt length of $n_{max\_len}=20$. For the EMAIL dataset, we maintained similar settings, adjusting only the maximum prompt length to $n_{max\_len}=100$. For the GPT-J-6B model, the following hyperparameters were used: on the SSN dataset, we targeted layers $3$-$8$ with a loss break of $0.1$, loss prediction probability coefficient of $20$, learning rate of $0.05$, $25$ gradient steps, and a maximum prompt length of $n_{max\_len}=200$. The EMAIL dataset configuration was similar, with a reduced maximum prompt length of $n_{max\_len}=50$. Other parameters remained consistent with the original paper’s recommendations. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.5 Head-Projection Baseline
We followed the original implementation with a modified objective to prevent the target from appearing in the top $k=100$ logit distribution elements. For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration used a loss break of $1.9\times 10^{-3}$, loss prediction probability coefficient of $3.08\times 10^{-2}$, learning rate of $1.03\times 10^{-2}$, and $10$ gradient steps. The EMAIL used a loss break of $7.6\times 10^{-4}$, loss prediction probability coefficient of $2.58\times 10^{-1}$, learning rate of $9.19\times 10^{-3}$, and $25$ gradient steps. For the GPT-J-6B model we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset we used a loss break of $6.596\times 10^{-2}$, loss prediction probability coefficient of $3.271$, learning rate of $2.045\times 10^{-1}$, and $25$ gradient steps. The EMAIL dataset configuration used similar layer targeting but modified the hyperparameters, with a loss break of $6\times 10^{-4}$, loss prediction probability coefficient of $6.733\times 10^{-1}$, learning rate of $1.868\times 10^{-1}$, and $5$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.6 Max-Entropy Baseline
For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration employed a loss break of $1.99\times 10^{-2}$, loss prediction probability coefficient of $20.5$, learning rate of $7.6\times 10^{-3}$, and $25$ gradient steps. The EMAIL dataset adjusted these parameters, using a loss break of $1.75\times 10^{-3}$, loss prediction probability coefficient of $1.35\times 10^{-2}$, learning rate of $9.2\times 10^{-3}$, and $5$ gradient steps. For the GPT-J-6B model, we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset, we used a loss break of $7.0559\times 10^{-2}$, loss prediction probability coefficient of $25.9$, learning rate of $1.937\times 10^{-1}$, and $5$ gradient steps. The EMAIL dataset used a loss break of $3.634\times 10^{-2}$, loss prediction probability coefficient of $1.907\times 10^{-2}$, learning rate of $6.65\times 10^{-2}$, and $10$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ E.1 Details and Examples
All methods in this work unlearn the same target tokens from sequences containing sensitive information. Our approach strategically selects the rarest tokens $T\subseteq S$ within each sequence $S$ to minimize unnecessary model perturbation while effectively isolating sensitive content. Token rarity is approximated using token ID assignments, where higher ID values indicate rarer tokens.
In our experiments, we unlearn two tokens ($|T|=2$) per target sequence. For example, consider the email address lewis.david@email.com, tokenized as [le, wis, .d, avid, @email, .com] with token IDs [273, 49143, 962, 15567, 72876, 916]. Here, we select wis and avid as the rarest tokens while excluding common domain-related tokens like @email.com.
For the URL dataset, we exclude common substrings such as [http, https, www, ://, /, ", -] from unlearning, as they do not encapsulate meaningful information and remain easily inferable. Similarly, in the SSN dataset, we exclude [-] for the same reason.
This strategy ensures the removal of the most unique, information-dense tokens that are most likely to contain sensitive identifiers while minimizing unnecessary model perturbation.